# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 2.3 Improved Convergence with Additional Structures
A desirable geometric structure that can facilitate the convergence of IPGD is the so-called strict saddle property on the implicit region $\mathcal{M}$, which we introduce next.
@@
We say the function $f$ satisfies the $(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}})$$-\mathcal{M}$-strict saddle property for some parameters $\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}}>0$ if, for any $\mathbf{x}\in\mathcal{M}$, at least one of following conditions holds:
@@
Intuitively, the $\mathcal{M}$-strict saddle property partitions the implicit region $\mathcal{M}$ into three distinct areas: $(i)$ a region where the gradient norm is large; $(ii)$ a region containing only approximate strict saddle points, i.e., points with sufficiently negative curvature; and $(iii)$ a neighborhood surrounding the exact $\mathcal{M}$-SOSPs.
This property is akin to the global strict saddle property [42, Assumption A2], with the key distinction that it is only imposed on the implicit region $\mathcal{M}$, whereas the global variant must hold for every $\mathbf{x}\in\mathbb{R}^{d}$. While the global strict saddle property holds for generic functions [18, Theorem 2.9], the associated parameters $(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}})$ tend to degrade as the level of over-parameterization increases, leading to overly conservative convergence rate for the algorithm. Therefore, imposing this condition solely on the implicit region $\mathcal{M}$—rather than the entire $\mathbb{R}^{d}$—can lead to less conservative parameter choices. For example, several problems, including matrix sensing, matrix completion, and tensor decomposition, have been shown to exhibit strict saddle properties with desirable parameters within their implicit regions [42, 19]. [^4] The original results consider a slightly different setting, assuming the problem is exactly-parameterized, whereas we address the over-parameterized setting, where $\mathbf{x}\in\mathcal{M}\subset\mathbb{R}^{d}$ with $\dim(\mathcal{M})=k<d$. However, a simple lifting trick can establish the equivalence between these cases. In contrast, in over-parameterized regimes, the global strict saddle property is conjectured to hold only with unfavorable parameters, resulting in poor convergence rates [50, Page 5].
In conjunction with Theorem 5, if the function $f$ additionally satisfies the $(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}})$$-\mathcal{M}$-strict saddle property, then IPGD converges to an $\bar{\epsilon}_{\mathcal{M}}$-neighborhood of an exact $\mathcal{M}$-SOSP, while staying close to the implicit region $\mathcal{M}$.
@@
Additionally, Proposition 6 offers another distinct advantage: under the assumption that $\bar{\epsilon}=\Theta(1)$, the required bound on the deviation rate simplifies to $R(\tau)=\tilde{O}\left(1/\Delta_{f}\right)$, which is a significant relaxation compared to the stricter requirement $R(\tau)=\tilde{O}\left(\epsilon^{2}/\Delta_{f}\right)$ in Theorem 5. As we will show later, both $\bar{\epsilon}=\Theta(1)$ and $R(\tau)=\tilde{O}\left(1/\Delta_{f}\right)$ can be established for the over-parameterized matrix sensing.
@@
The $(\alpha,\beta,\zeta)$$-\mathcal{M}$-regularity property is a weaker condition than local strong convexity. In fact, it can be shown that gradient Lipschitzness and one-point strong convexity together imply this property (see [49, Section II.B] for a proof). However, unlike local strong convexity—which typically fails to hold even in the exactly-parameterized regime [^5] Strong convexity provably fails within any neighborhood if the solution set forms a connected set. This occurs in various models, including $s$-sparse recovery and low-rank matrix recovery.—the $(\alpha,\beta,\zeta)$$-\mathcal{M}$-regularity property is generally easier to establish. For example, it has been verified for several nonconvex optimization problems, such as matrix factorization [20], matrix sensing [41], and matrix completion [51] in the exactly-parameterized regime.
@@
Note that the improvement in the iteration count—from $\operatorname{poly}(1/\epsilon)$ to $\log(1/\epsilon)$—enabled by the regularity condition naturally relaxes the dependence of the deviation rate $R(\tau)$ on $\epsilon$ to $\log^{-1}(1/\epsilon)$. This is significantly less restrictive than the requirement in Theorem 5, where $R(\tau)$ must scale as $\epsilon^{2}$.
@@
Algorithm 2 IPGD with local improvement: IPGD+$(\gamma,\epsilon,\eta,\mathbf{x}_{0},\rho,\Delta_{f},\chi,T^{\prime})$
1:$\mathbf{x}_{0}\leftarrow\textsc{IPGD}(\gamma,\epsilon,\eta,\mathbf{x}_{0},\rho,\Delta_{f},\chi)$
@@
Theorem 8 (Global near-linear convergence of IPGD+ under strict saddle and regularity properties).
Suppose that $f$ is ${(L,\mathcal{M},\tau)}$-gradient-Lipschitz and ${(\rho,\mathcal{M},\tau)}$-Hessian-Lipschitz with $\tau\geq\epsilon/L$. Moreover, suppose that $f$ satisfies the $(\alpha,\beta,\zeta)$$-\mathcal{M}$-regularity property and $(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}})$$-\mathcal{M}$-strict saddle property with $\bar{\epsilon}_{\mathcal{M}}\leq\zeta$. Suppose that the initial point satisfies $\left\lVert\bm{\mathbf{x}}^{\perp}_{0}\right\rVert=O\left(\min\left\{\eta\alpha\cdot\frac{\epsilon}{L},\frac{1}{\eta\rho T}\right\}\right)$. Furthermore, suppose that Conditions C1 and C3 from Theorem 5 hold. Let $\bar{\epsilon}=O\left(\min\left\{\bar{\epsilon}_{\mathrm{g}},\frac{\bar{\epsilon}_{\mathrm{H}}^{2}}{\rho}\right\}\right)$ and $T^{\prime}=O\left(\frac{1}{\eta\alpha}{\log\left(\frac{L\zeta}{\epsilon}\right)}\right)$. Then, with probability of at least $1-\chi$, IPGD+ (Algorithm 2) with any perturbation radius satisfying $\gamma=O\Big{(}\min\Big{\{}\frac{\bar{\epsilon}^{3/2}\cdot\epsilon}{\sqrt{\rho}L\Delta_{f}},\frac{\bar{\epsilon}^{7/2}}{\rho^{3/2}\Delta_{f}^{2}}\Big{\}}\cdot\log^{-7}\left(\frac{\rho d\Delta_{f}}{\chi\bar{\epsilon}}\right)\Big{)}$ and stepsize $\eta\leq\frac{1}{10\max\{\alpha,\beta,L\}}$ outputs a solution satisfying
@@
We note that IPGD+ is similar to the PGD with local improvement (referred to as PGDli) introduced in [20]. However, IPGD+ has a slower convergence rate by a factor of $\log^{3}(1/\epsilon)$. This slowdown stems from the use of a smaller perturbation radius, which is necessary to enable implicit regularization and ensure that the iterates remain close to the implicit region $\mathcal{M}$.
@@ 3 Establishing Convergence of IPGD to an Approximate SOSP
To establish the convergence of IPGD, our analysis carefully refines the convergence analysis of PGD from [20, 23], extending it to accommodate infinitesimal perturbations.
To establish the convergence of IPGD to an approximate SOSP, we divide our analysis into two parts:
  • Bounding the iteration count: We derive an upper bound on the maximum number of iterations that IPGD can take before meeting its termination criterion, thereby characterizing its convergence rate.
@@
By combining these observations with the definitions of $N_{\mathrm{large}}$ and $N_{\mathrm{perturb}}$, we obtain $T\leq N_{\mathrm{large}}+N_{\mathrm{perturb}}\cdot T_{\mathrm{escape}}$, thereby completing the proof. ∎
Based on the above lemma, in order to bound the iteration count, it suffices to provide separate bounds on $N_{\mathrm{large}}$ and $N_{\mathrm{perturb}}$. Towards this goal, we utlize a classic descent lemma for GD updates.
Lemma 10 (Descent lemma for GD updates; [52, Section 1.2.3]).
@@
Suppose that the conditions of Theorem 2 are satisfied. Then, IPGD terminates within $T\leq\frac{4C^{2}\Delta_{f}}{\eta\epsilon^{2}}\left(\log^{4}\left(\frac{1}{\gamma}\right)+\log^{4}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)$ iterations.
@@
$f(\mathbf{x}_{t+1})-f(\mathbf{x}_{t})\leq-\frac{1}{2}\eta\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert^{2}\leq-\frac{1}{2}\eta G=-\frac{1}{2}\eta\left(\frac{1}{C}\cdot\frac{\epsilon}{\log^{2}(1/\gamma)}\right)=-\frac{\eta\epsilon}{2C\log^{2}(1/\gamma)}$ (20)
@@
$N_{\mathrm{large}}\leq\frac{\Delta_{f}}{\frac{\eta\epsilon}{2C\log^{2}(1/\gamma)}}=\frac{2C\Delta_{f}\log^{2}(1/\gamma)}{\eta\epsilon}.$ (21)
Next, we provide an upper bound on $N_{\mathrm{perturb}}$. Let $t_{k}$ denote the iteration in which the $k$-th perturbation is applied to $\mathbf{x}_{t_{k}}$. For every $11\leq k\leq N_{\mathrm{perturb}}-$ 1, we must have
@@
$\leq\frac{2C\Delta_{f}\log^{2}(1/\gamma)}{\eta\epsilon}+\frac{C\sqrt{\rho}\Delta_{f}\left(\log^{3}\left(\frac{1}{\gamma}\right)+\log^{3}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)}{\epsilon^{3/2}}\cdot\frac{C\left(\log\left(\frac{1}{\gamma}\right)+\log\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)}{\eta\sqrt{\rho\epsilon}}$
@@
Next, we establish that, upon termination, the returned solution is indeed an approximate SOSP. To this goal, we rely on the following key proposition, which encapsulates the main novelty of our proof.
@@
Now, we proceed with the proof of Proposition 12. We first introduce the following lemma from [23].
Lemma 13 (Improve or localize, adapted from Lemma 5.4 in [23]).
@@
Our next result is at the core of the guaranteed escape of IPGD from an approximate strict saddle point. To explain the intuition behind this result, let $\mathbf{z}$ be an approximate strict saddle point, satisfying $\left\lVert\nabla f(\mathbf{z})\right\rVert\leq\epsilon$ and $\lambda_{\min}(\nabla^{2}f(\mathbf{z}))<-\sqrt{\rho\epsilon}$. Let $\mathbf{v}$ be the eigenvector corresponding to the most negative eigenvalue of $\nabla^{2}f(\mathbf{z})$. Consider any two initial points $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}$ such that $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}\in\mathbb{B}_{\mathbf{z}}(\gamma)$ and $\mathbf{x}_{0}-\mathbf{x}_{0}^{\prime}=\omega\cdot\mathbf{v}$ for some $\omega>0$. Our next result demonstrates that GD, when initialized at one of these points, must decrease the objective function rapidly, enabling an escape from $\mathbf{z}$.
@@
Let $\mathbf{z}$ be an approximate strict saddle point satisfying $\left\lVert\nabla f(\mathbf{z})\right\rVert\leq\epsilon$ and $\lambda_{\min}(\nabla^{2}f(\mathbf{z}))<-\sqrt{\rho\epsilon}$. Let $\mathbf{v}$ be the eigenvector corresponding to the most negative eigenvalue of $\nabla^{2}f(\mathbf{z})$. Consider two points $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}$ satisfying $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}\in\mathbb{B}_{\mathbf{z}}(\gamma)$ and $\mathbf{x}_{0}-\mathbf{x}_{0}^{\prime}=\omega\cdot\mathbf{v}$ for some $0<\omega\leq 2\gamma$. Then, the two sequences $\{\mathbf{x}_{t}\}_{t=0}^{T}$ and $\{\mathbf{x}_{t}^{\prime}\}_{t=0}^{T}$ generated by GD updates with constant stepsize $\eta\leq\frac{1}{L}$ satisfy the following property within $T\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$ iterations:
@@
Let $T$ be the first iteration satisfying $\max\{\left\lVert\mathbf{x}_{T}-\mathbf{z}\right\rVert,\left\lVert\mathbf{x}_{T}^{\prime}-\mathbf{z}\right\rVert\}\geq\phi$. First, we show that $T\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$. Evidently, we have $\max\{\left\lVert\mathbf{x}_{t}-\mathbf{z}\right\rVert,\left\lVert\mathbf{x}_{t}^{\prime}-\mathbf{z}\right\rVert\}<\phi$ for all $10\leq t\leq T-$ 1. We denote $\bm{\delta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\prime}$, $\mathbf{H}=\nabla^{2}f(\mathbf{z})$, and $\bm{\Delta}_{t}=\int_{0}^{1}\left(\nabla^{2}f\left(\mathbf{x}_{t}^{\prime}+s\bm{\delta}_{t}\right)-\mathbf{H}\right)\mathrm{d}s$. Then, for every $10\leq t\leq T-$ 1, we have
@@
$\bm{\delta}_{t}$ $=\bm{\delta}_{t-1}-\eta\left(\nabla f\left(\mathbf{x}_{t-1}\right)-\nabla f\left(\mathbf{x}_{t-1}^{\prime}\right)\right)$ (30)
$=(\mathbf{I}-\eta\mathbf{H})\bm{\delta}_{t-1}-\eta\bm{\Delta}_{t-1}\bm{\delta}_{t-1}$
@@
Upon defining $p_{t}:=\left\lVert(\mathbf{I}-\eta\mathbf{H})^{t}\bm{\delta}_{0}\right\rVert$ and $q_{t}:=\left\lVert\eta\sum_{s=0}^{t-1}(\mathbf{I}-\eta\mathbf{H})^{t-s-1}\bm{\Delta}_{s}\bm{\delta}_{s}\right\rVert$, the above equality implies that

### v2

@@ 2.3 Improved Convergence with Additional Structures
A desirable geometric structure that can facilitate a faster convergence of IPGD is the so-called strict saddle property on the implicit region $\mathcal{M}$, which we introduce next.
@@
We say the function $f$ satisfies the $(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}})$-$\mathcal{M}$-strict saddle property for some parameters $\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}}>0$ if, for any $\mathbf{x}\in\mathcal{M}$, at least one of the following conditions holds:
@@
Intuitively, the $\mathcal{M}$-strict saddle property partitions the implicit region $\mathcal{M}$ into three areas: (i) regions with large gradient norms; (ii) regions containing approximate SSPs with sufficiently negative curvature; and (iii) neighborhoods around exact $\mathcal{M}$-SOSPs.
This property is akin to the global strict saddle property [ge2017no, Assumption A2], with the key distinction that it is only imposed on the implicit region $\mathcal{M}$, whereas the global variant must hold for every $\mathbf{x}\in\mathbb{R}^{d}$. While the global strict saddle property holds for generic functions [davis2019active, Theorem 2.9], the associated parameters $(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}})$ tend to degrade as the level of over-parameterization increases, leading to overly conservative convergence rate for the algorithm. Therefore, imposing this condition solely on the implicit region $\mathcal{M}$—rather than the entire $\mathbb{R}^{d}$—can lead to less conservative parameter choices. For example, several problems, including matrix sensing, matrix completion, and tensor decomposition, have been shown to exhibit strict saddle properties with desirable parameters within their implicit regions [ge2017no, ge2015escaping]. [^2] The original results consider a slightly different setting, assuming the problem is exactly-parameterized, whereas we address the over-parameterized setting, where $\mathbf{x}\in\mathcal{M}\subset\mathbb{R}^{d}$ with $\dim(\mathcal{M})=k<d$. However, a simple lifting trick can establish the equivalence between these cases. In contrast, in over-parameterized regimes, the global strict saddle property is conjectured to hold only with unfavorable parameters, resulting in poor convergence rates [zhang2021sharp, Page 5].
We note that, in general, the iterates of IPGD do not necessarily belong to the implicit region $\mathcal{M}$. Consequently, it is not immediately clear why the $\mathcal{M}$-strict saddle property—a property defined solely on $\mathcal{M}$—should facilitate the convergence of iterates that lie outside this region. Nonetheless, our next proposition shows that the strict saddle property, when restricted to $\mathcal{M}$, is sufficient to improve the convergence of IPGD.
@@
Additionally, Proposition 2 offers another distinct advantage: under the assumption that $\bar{\epsilon}=\Theta(1)$, the required bound on the deviation rate simplifies to $R(\tau)=\tilde{O}\left(1/\Delta_{f}\right)$, which is a significant relaxation compared to the stricter requirement $R(\tau)=\tilde{O}\left(\epsilon^{2}/\Delta_{f}\right)$ in Theorem 3. As we will show later, both $\bar{\epsilon}=\Theta(1)$ and $R(\tau)=\tilde{O}\left(1/\Delta_{f}\right)$ can be provably established for the over-parameterized matrix sensing.
@@
The $(\alpha,\beta,\zeta)$-$\mathcal{M}$-regularity property is a weaker condition than local strong convexity. In fact, it can be shown that gradient Lipschitzness and one-point strong convexity together imply this property (see [chi2019nonconvex, Section II.B] for a proof). However, unlike local strong convexity—which typically fails to hold even in the exactly-parameterized regime [^3] Strong convexity provably fails within any neighborhood if the solution set forms a connected set. This occurs in various models, including sparse recovery and low-rank matrix recovery.—the $(\alpha,\beta,\zeta)$-$\mathcal{M}$-regularity property is generally easier to establish. For example, it has been verified for several nonconvex optimization problems, such as matrix factorization [jin2017escape], matrix sensing [bhojanapalli2016global], and matrix completion [zheng2016convergence] in the exactly-parameterized regime.
@@
Note that the improvement in the iteration count—from $\operatorname{poly}(1/\epsilon)$ to $\log(1/\epsilon)$—enabled by the regularity condition relaxes the dependence of the deviation rate $R(\tau)$ on $\epsilon$ to $\log^{-1}(1/\epsilon)$. This is significantly less restrictive than the requirement in Theorem 3, where $R(\tau)$ must scale as $\epsilon^{2}$.
@@
[…]
@@
Theorem 4 (Global near-linear convergence of IPGD+).
Suppose that $f$ is $(L,\mathcal{M},\tau)$}-gradient-Lipschitz and $(\rho,\mathcal{M},\tau)$}-Hessian-Lipschitz with $\tau\geq\epsilon/L$. Suppose that $f$ satisfies the $(\alpha,\beta,\zeta)$-$\mathcal{M}$-regularity property and $(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}})$-$\mathcal{M}$-strict saddle property with $\bar{\epsilon}_{\mathcal{M}}\leq\zeta$. Suppose that the initial point satisfies $\left\lVert\bm{\mathbf{x}}^{\perp}_{0}\right\rVert=O\left(\min\left\{\eta\alpha\cdot\frac{\epsilon}{L},\frac{1}{\eta\rho T}\right\}\right)$. Suppose that Conditions C1 and C3 from Theorem 3 hold. Let $\bar{\epsilon}=O\left(\min\left\{\bar{\epsilon}_{\mathrm{g}},\frac{\bar{\epsilon}_{\mathrm{H}}^{2}}{\rho}\right\}\right)$ and $T^{\prime}=O\left(\frac{1}{\eta\alpha}{\log\left(\frac{L\zeta}{\epsilon}\right)}\right)$. Then, with probability of at least $1-\chi$, IPGD+ (Algorithm 2) with any perturbation radius $\gamma=O\Big(\min\Big\{\frac{\bar{\epsilon}^{3/2}\cdot\epsilon}{\sqrt{\rho}L\Delta_{f}},\frac{\bar{\epsilon}^{7/2}}{\rho^{3/2}\Delta_{f}^{2}}\Big\}\cdot\log^{-7}\left(\frac{\rho d\Delta_{f}}{\chi\bar{\epsilon}}\right)\Big)$ and stepsize $\eta\leq\frac{1}{10\max\{\alpha,\beta,L\}}$ outputs a solution satisfying
@@
We note that IPGD+ is similar to the PGD with local improvement (referred to as PGDli) introduced in [jin2017escape]. However, IPGD+ has a slower convergence rate by a factor of $\log^{3}(1/\epsilon)$. This slowdown stems from the use of a smaller perturbation radius, which is necessary to enable implicit regularization and ensure that the iterates remain close to the implicit region $\mathcal{M}$.
@@ 3 Establishing Convergence of IPGD to an Approximate SOSP
To prove the convergence of IPGD, we carefully refine the analysis of PGD from [jin2017escape, jin2021nonconvex], extending it to handle infinitesimal perturbations. Our analysis proceeds in two parts:
  • Bounding the iteration count: We derive an upper bound on the number of iterations that IPGD can take before meeting its termination criterion, characterizing its convergence rate.
@@
By combining these observations with the definitions of $N_{\mathrm{large}}$ and $N_{\mathrm{perturb}}$, we obtain $T\leq N_{\mathrm{large}}+N_{\mathrm{perturb}}\cdot T_{\mathrm{escape}}$, thereby completing the f. ∎
Based on the above lemma, to bound the iteration count, it suffices to provide separate bounds on $N_{\mathrm{large}}$ and $N_{\mathrm{perturb}}$. Towards this goal, we utilize a classic descent lemma for GD updates.
Lemma 2 (Descent lemma for GD updates; [nesterov2018lectures, Section 1.2.3]).
@@
Under the conditions of Theorem 2, IPGD terminates within $T\leq\frac{4C^{2}\Delta_{f}}{\eta\epsilon^{2}}\left(\log^{4}\left(\frac{1}{\gamma}\right)+\log^{4}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)$ iterations.
@@
$f(\mathbf{x}_{t+1})-f(\mathbf{x}_{t})\leq-\frac{1}{2}\eta\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert^{2}\leq-\frac{1}{2}\eta G^{2}=-\frac{1}{2}\eta\left(\frac{1}{C^{2}}\cdot\frac{\epsilon^{2}}{\log^{4}(1/\gamma)}\right)=-\frac{\eta\epsilon^{2}}{2C^{2}\log^{4}(1/\gamma)}$ (21)
@@
$N_{\mathrm{large}}\leq\frac{\Delta_{f}}{\frac{\eta\epsilon^{2}}{2C^{2}\log^{4}(1/\gamma)}}=\frac{2C^{2}\Delta_{f}\log^{4}(1/\gamma)}{\eta\epsilon^{2}}.$ (22)
Next, we provide an upper bound on $N_{\mathrm{perturb}}$. Let $t_{k}$ denote the iteration in which the $k$-th perturbation is applied to $\mathbf{x}_{t_{k}}$. For every $1\leq k\leq N_{\mathrm{perturb}}-1$, we must have
@@
$\leq\frac{2C^{2}\Delta_{f}\log^{4}(1/\gamma)}{\eta\epsilon^{2}}+\frac{C\sqrt{\rho}\Delta_{f}\left(\log^{3}\left(\frac{1}{\gamma}\right)+\log^{3}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)}{\epsilon^{3/2}}\cdot\frac{C\left(\log\left(\frac{1}{\gamma}\right)+\log\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)}{\eta\sqrt{\rho\epsilon}}$
@@
Next, we establish that the returned solution is indeed an approximate SOSP. To this goal, we rely on the following key proposition, which encapsulates the main novelty of our proof.
@@
Now, we proceed with the proof of Proposition 5. We first introduce the following lemma.
Lemma 3 (Improve or localize, adapted from Lemma 5.4 in [jin2021nonconvex]).
@@
Our next result is at the core of the guaranteed escape of IPGD from an approximate SSP. To explain the intuition behind this result, let $\mathbf{z}$ be an approximate SSP, satisfying $\left\lVert\nabla f(\mathbf{z})\right\rVert\leq\epsilon$ and $\lambda_{\min}(\nabla^{2}f(\mathbf{z}))<-\sqrt{\rho\epsilon}$. Let $\mathbf{v}$ be the eigenvector corresponding to the most negative eigenvalue of $\nabla^{2}f(\mathbf{z})$. Consider any two initial points $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}$ such that $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}\in\mathbb{B}_{\mathbf{z}}(\gamma)$ and $\mathbf{x}_{0}-\mathbf{x}_{0}^{\prime}=\omega\cdot\mathbf{v}$ for some $\omega>0$. Our next result demonstrates that GD, when initialized at one of these points, must decrease the objective function rapidly, enabling an escape from $\mathbf{z}$.
@@
Let $\mathbf{z}$ be an approximate SSP satisfying $\left\lVert\nabla f(\mathbf{z})\right\rVert\leq\epsilon$ and $\lambda_{\min}(\nabla^{2}f(\mathbf{z}))<-\sqrt{\rho\epsilon}$. Let $\mathbf{v}$ be the eigenvector corresponding to the most negative eigenvalue of $\nabla^{2}f(\mathbf{z})$. Consider two points $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}$ satisfying $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}\in\mathbb{B}_{\mathbf{z}}(\gamma)$ and $\mathbf{x}_{0}-\mathbf{x}_{0}^{\prime}=\omega\cdot\mathbf{v}$ for some $0<\omega\leq 2\gamma$. Then, the two sequences $\{\mathbf{x}_{t}\}_{t=0}^{T}$ and $\{\mathbf{x}_{t}^{\prime}\}_{t=0}^{T}$ generated by GD updates with constant stepsize $\eta\leq\frac{1}{L}$ satisfy the following property within $T\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$ iterations:
@@
Let $T$ be the first iteration satisfying $\max\{\left\lVert\mathbf{x}_{T}-\mathbf{z}\right\rVert,\left\lVert\mathbf{x}_{T}^{\prime}-\mathbf{z}\right\rVert\}\geq\phi$. First, we show that $T\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$. Evidently, we have $\max\{\left\lVert\mathbf{x}_{t}-\mathbf{z}\right\rVert,\left\lVert\mathbf{x}_{t}^{\prime}-\mathbf{z}\right\rVert\}<\phi$ for all $0\leq t\leq T-1$. We denote $\bm{\delta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\prime}$, $\mathbf{H}=\nabla^{2}f(\mathbf{z})$, and $\bm{\Delta}_{t}=\int_{0}^{1}\left(\nabla^{2}f\left(\mathbf{x}_{t}^{\prime}+s\bm{\delta}_{t}\right)-\mathbf{H}\right)\mathrm{d}s$. Then, for every $0\leq t\leq T-1$:
@@
$\bm{\delta}_{t}$ $=\bm{\delta}_{t-1}-\eta\left(\nabla f\left(\mathbf{x}_{t-1}\right)-\nabla f\left(\mathbf{x}_{t-1}^{\prime}\right)\right)=(\mathbf{I}-\eta\mathbf{H})\bm{\delta}_{t-1}-\eta\bm{\Delta}_{t-1}\bm{\delta}_{t-1}$ (31)
@@
Upon defining $p_{t}:=\left\lVert(\mathbf{I}-\eta\mathbf{H})^{t}\bm{\delta}_{0}\right\rVert$ and $q_{t}:=\left\lVert\eta\sum_{s=0}^{t-1}(\mathbf{I}-\eta\mathbf{H})^{t-s-1}\bm{\Delta}_{s}\bm{\delta}_{s}\right\rVert$, we have